# Bridging AI and Science: Implications from a Large-Scale Literature Analysis of AI4Science

### v1

@@ Abstract.
Artificial Intelligence has proven to be a transformative tool for advancing scientific research across a wide range of disciplines. However, a significant gap still exists between AI and scientific communities, limiting the full potential of AI methods in driving broad scientific discovery. Existing efforts in bridging this gap have often relied on qualitative examination of small samples of literature, offering a limited perspective on the broader AI4Science landscape. In this work, we present a large-scale analysis of the AI4Science literature, starting by using large language models to identify scientific problems and AI methods in publications from top science and AI venues. Leveraging this new dataset, we quantitatively highlight key disparities between AI methods and scientific problems in this integrated space, revealing substantial opportunities for deeper AI integration across scientific disciplines. Furthermore, we explore the potential and challenges of facilitating collaboration between AI and scientific communities through the lens of link prediction. Our findings and tools aim to promote more impactful interdisciplinary collaborations and accelerate scientific discovery through deeper and broader AI integration.
@@ 1. Introduction
The 2024 Nobel Prizes in Physics and Chemistry were both awarded to Artificial Intelligence (AI) researchers. In particular, the developers of AlphaFold (Jumper et al., 2021) were recognized for their groundbreaking work, which is based on the Transformer network and has revolutionized the prediction of protein structures, providing a transformative approach to solving complex biological problems. This is just one example of how AI has emerged as a powerful tool to facilitate scientific research across various disciplines (Wang et al., 2023; Leontidis, 2024; Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
@@
However, despite the transformative potential of using AI for advancing scientific research (AI4Science), a considerable gap persists between AI and scientific communities, hindering the full exploitation of AI for scientific discovery. On one side, advanced AI methodologies might remain mysterious or underutilized by scientists; on the other side, AI researchers may lack awareness of the specific challenges and potential applications in scientific domains, missing the opportunities for interdisciplinary collaborations.
@@
Existing efforts to identify and bridge this gap have predominantly involved qualitative reviews of AI’s application in science (Wang et al., 2023; Leontidis, 2024), particularly in specialized areas like drug discovery and materials science (Mak et al., 2023; Guo et al., 2021; Xu et al., 2024). Based on qualitative examination of small samples of literature, these reviews often rely on heuristic insights from domain experts to suggest potential uses of AI in solving scientific problems. While valuable, such focused reviews are limited in providing a comprehensive and diverse perspective on the AI4Science landscape.
@@
While recent research from the Science of Science community has conducted larger-scale studies that quantify AI’s impact on publications in scientific fields (Gao and Wang, 2023; Duede et al., 2024), they often heavily rely on established scientific taxonomies in “AI-heavy” fields. There remains a lack of delivering a holistic, dynamic, and data-driven overview of AI4Science. Such analysis is crucial for understanding the barriers and identifying new opportunities for deep engagement of fast-evolving AI methodology in scientific research.
@@
In this work, we aim to depict the AI4Science landscape via a large-scale and comprehensive analysis of relevant literature in both scientific and AI communities over the past decade. We start by using large language models (LLMs) to identify the scientific problems and AI methods addressed in publications from top science and AI venues, through which we assemble a novel and balanced AI4Science dataset to analyze the role of AI in scientific research (Sec. 3). Then with projection maps (Fig. 1) and a bipartite graph (Fig. 2) constructed from this dataset, we quantitatively reveal the discrepancies between AI and science, leading to several key findings and implications (Sec. 4). Finally, through the lens of link prediction, we demonstrate the potential and challenges in advancing the bridge between AI and science (Sec. 5). We anticipate that our findings and tools will foster more effective interdisciplinary collaborations, ultimately accelerating scientific discovery through a more diverse and deeper integration of AI methods.
Figure 1. Illustration of the AI4Science dataset: (1) LLM-based extraction of scientific problem $p_{i}$, AI method $m_{i}$, and AI usage $u_{i}$ from an example publication. (2) 2D projections of the extracted scientific problems $\{p_{i}\}$ and AI methods $\{m_{i}\}$. Each dot represents a publication, with coordinates obtained by reducing the dimensionality of the problem/method embeddings. (a) Green dots correspond to AI4Science work, where AI is utilized to address scientific problems; Orange dots are papers addressing scientific problems with non-AI techniques; (b) Purple dots indicate using AI to address non-scientific problems. The visualizations reveal a noticeable discrepancy in the distribution of AI4Science work (green) versus non-AI4Science work (orange and purple) in both the problem and method spaces.
Figure 2. The AI-Science bipartite graph and node distributions. (a) The bipartite graph with scientific problem clusters and AI method clusters as nodes. The size of each node corresponds to its unweighted degree, representing the number of AI methods applied to a scientific problem, or the number of scientific problems that applied an AI method. For clarity, we omit edges representing fewer than five publications, and isolated nodes are removed thereafter. The distribution of links indicates an imbalanced connectivity. (b-c) The degree of AI method nodes follows a log-normal distribution, and the degree distribution of scientific problem nodes is even more heavily tailed, suggesting the existence of “hubs” in linking AI and science.
Notation Value
Publication year range 2014–2024
# of publications $N$ 159,295
# of scientific problem extractions |$|\{p_{i}\}|$ 123,773
# of scientific problem clusters |$|\{P_{i}\}|$ 369
# of AI method extractions |$|\{m_{i}\}|$ 35,522
# of AI method clusters |$|\{M_{i}\}|$ 345
# of AI usage extractions |$|\{u_{i}\}|$ 35,522
# of AI4Science publications 7,024
2. Related Work
2.1. Qualitative Literature Reviews of AI4Science
While these reviews provide valuable insights, they often rely on interdisciplinary experts to manually identify and categorize scientific problems and potential AI methods, limiting their scope. Such an approach often could not offer a comprehensive and data-driven overview of the broader “AI4Science” landscape.
2.2. Quantitative Literature Analysis
In addition to heuristic and qualitative literature reviews, a few recent papers from the Science of Science community have focused on quantitatively analyzing the AI4Science publications. For instance, Gao and Wang (2023) examined how AI benefits scientific research, finding that AI use is prevalent across various scientific disciplines, and that papers incorporating AI tend to receive a citation advantage. Similarly, Duede et al. (2024) explored AI’s engagement in different fields, uncovering an “oil-and-water” phenomenon – while AI-engaged work is spreading across disciplines, it does not integrate well with non-AI-engaged work within each specific field.
Although these studies offer valuable insights into AI4Science, they primarily focus on areas where AI and science are already “blended” effectively, and overlook the broader and underexplored parts of the landscape. As a result, these approaches often fail to offer insights into the disparities between AI and scientific communities. In contrast, our work offers a more comprehensive view of the AI4Science landscape by showing and quantifying the connections and gaps between scientific problems and AI methods, as well as insights to uncover potential collaborations.
3. A New Dataset of AI4Science
@@
A major challenge in understanding the disparities and potential of AI4Science is the lack of a comprehensive, balanced dataset. Existing literature data are often biased toward success stories, such as areas where AI and science are already deeply intertwined. Or they reflect the perspective of either the scientific or AI communities, but not both. For example, the AI methods discussed in the scientific literature may not represent the full range of advancements in AI, while the scientific applications explored in AI literature may fail to capture the nuances and challenges of real-world scientific problems. We take the initiative to create a broad and balanced literature dataset to provide a wide-angled lens for understanding AI4Science research.
@@ 3.1. Curation of AI4Science Literature
  • For science domains, we include three top multidisciplinary journals that represent cutting-edge and high-quality research (Nature, Science, and PNAS) and two of their subjournals (Nature Communications, and Science Advances).
  • For AI communities, we include six top conferences from the list of CSRankings.org, comprising two AI-focused venues (AAAI, IJCAI), three machine learning venues (ICLR, ICML, NeurIPS), and the Web Conference (WWW) as a venue of applied AI research.
Our data spans the past decade, from 2014 to 2024, including a total of 159,295 publications from both AI and science venues. More details of the data curation process can be found in Appendix A.1.
@@ 3.2. Extracting Scientific Problems and AI Methods with LLMs
To understand how AI is applied to science, we need to extract paired entries of scientific problems and AI solutions from relevant publications. Traditionally, this is accomplished through literature reviews and surveys which involve manual examinations of small samples of papers (Wang et al., 2023; Leontidis, 2024), a labor-intensive and time-consuming process. Recent quantitative studies in the Science of Science field have started using metadata entries, such as keywords, to identify AI usage and categorize scientific topics (Gao and Wang, 2023; Duede et al., 2024). While this method enables larger-scale analysis, it offers only a coarse-grained view, potentially missing the nuanced details of scientific problems and AI methods. Additionally, it can suffer from issues such as incomplete data entries, limiting the effectiveness of analysis.
@@
It is possible that a publication did not address a scientific problem or did not use an AI method, in which case the corresponding field(s) would remain empty. If a paper addresses a scientific problem using an AI method – i.e., $p_{i}$, $m_{i}$, and $u_{i}$ are all non-empty – it is considered as an AI4Science work. Figure 1 presents an example of the extraction results of an AI4Science publication.
@@
3.3. Semantic Clustering of Scientific Problems and AI methods
One of the key differences between this new AI4Science dataset and the data used in existing studies (Gao and Wang, 2023; Duede et al., 2024) is that our dataset contains detailed textual descriptions of the identified scientific problems and AI methods rather than pre-defined taxonomies, allowing for a deeper semantic analysis of their relationships.
To achieve this, we generate semantic embeddings for the scientific problems ${p_{i}}$, AI methods ${m_{i}}$, and AI usages ${u_{i}}$ using the InstructorEmbedding model (Su et al., 2022). We then apply LargeVis (Tang et al., 2016) to project these high-dimensional embeddings into 2D coordinates. Next, HDBSCAN (Campello et al., 2013) is utilized to group similar scientific problems and AI methods. For each cluster, we use GPT-4o to summarize the main topic based on frequent keywords and samples of extraction results. This process results in 369 and 345 clusters respectively, with corresponding cluster labels $P_{i}$ and $M_{i}$ to each data point. Fig. 1 displays the 2D visualizations, and Fig. 4-5 in the Appendix presents the maps with cluster names annotated. Details of the process can be found in Appendix A.3.
Unlike directly using keywords or categorizations from publication metadata (Gao and Wang, 2023; Duede et al., 2024), our approach employs a data-driven methodology. By using clustering and summarization, relevant AI methods discussed in both scientific and AI literature are grouped together, as are the scientific problems mentioned in both fields. This ensures greater consistency and generalizability of the data entries.
@@ 3.4. Dataset Overview and the Bipartite Graph
To better illustrate the connections between AI and science, we construct a bipartite graph based on the clusters. In this bipartite graph, scientific problem clusters and AI method clusters represent two types of nodes, and the publications act as edges connecting them. Formally, the graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{P_{i}\}\cup\{M_{i}\}$ and $\mathcal{E}=\mathcal{D}$. As shown in Fig. 2(a), the graph provides an intuitive visual representation of how AI is linked to scientific challenges and vice versa.
@@
Figure 3. Scientific problem clusters (a) and AI method clusters (b) are visualized as circles in the scatter plots. In both plots, the $x$-axis represents the total number of publications in each cluster, while the $y$-axis reflects the number of interdisciplinary AI4Science publications in each cluster. Three outstanding points are skipped for clarity1. The black lines show the regression results on the clusters. Clusters falling below the regression line highlight the under-investigated regions, where the integration of AI and science remains limited. These areas represent potential opportunities for further interdisciplinary collaboration.
@@ 4. Discrepancy Between AI and Science
Both Fig. 1 and 2 show a noticeable pattern of discrepancy between AI and science. This section quantitatively assesses these disparities, providing a deeper understanding of the gaps.
@@ 4.1. Uneven Distribution of AI4Science Research
The projection maps in Fig. 1 illustrate the distribution of AI4Science work within the semantic spaces of scientific problems and AI methods (the green dots among orange and purple dots). The uneven distributions suggest that AI4Science work is heavily clustered in certain subareas, while sit remains underrepresented in wide ranges of both the problem and the method spaces.
In Fig. 3, we plot the clusters of scientific problems and AI methods, showing the relation between cluster sizes and the presence of AI4Science publications within each cluster. The slopes of the regression lines represent the average proportions of AI4Science work in the literature of each scientific problem cluster or AI method cluster. Clusters below the regression line indicate problems or AI methods in which the integration of AI4Science is underrepresented (noted the under-investigated regions), suggesting potential opportunities for further interdisciplinary collaboration.
Under-investigated scientific problems.
In the bottom-right section of Fig. 3(a), below the regression line, several scientific problem clusters are highlighted as being rarely addressed by AI techniques. Notable examples include Magnetic Materials, Asymmetric Synthesis, Lithium-Ion Battery Challenges, Antibiotic Resistance, Stem Cell Differentiation, Viral Mechanisms, Climate Change Emissions, and Enzyme Catalysis. These scientific problems could benefit from more engagement of AI methods.
Under-investigated AI methods.
Similarly, Fig. 3(b) reveals AI techniques that have not yet been widely applied to sciences, including: Attention Mechanism, Gradient-Based Methods/Gradient Descent, Embedding Techniques, Policy Optimization, Regularization Techniques, Theoretical Analysis, Adversarial Training, Contrastive Learning, and Variational Inference. These techniques could have been explored in a wider range of scientific applications.
11footnotetext: Three outstanding data points are skipped from the scatter plots for visualization clarity: Neural Network Challenges (838, 653) in (a), Machine Learning (601, 446) and Deep Learning (577, 308) in (b).
Node type Partition Avg. Deg.
Scientific problem clusters (n=369) All nodes 15.7
Well-investigated 63.2
Under-investigated 6.3
AI method clusters (n=345) All nodes 16.5
Well-investigated 29.2
Under-investigated 8.5
Table 2. Average degree of nodes in the bipartite graph.

### v2

@@ Abstract.
Artificial Intelligence has proven to be a transformative tool for advancing scientific research across a wide range of disciplines. However, a significant gap still exists between AI and scientific communities, limiting the full potential of AI methods in driving broad scientific discovery. Existing efforts in identifying and bridging this gap have often relied on qualitative examination of small samples of literature, offering a limited perspective on the broader AI4Science landscape. In this work, we present a large-scale analysis of the AI4Science literature, starting by using large language models to identify scientific problems and AI methods in publications from top science and AI venues. Leveraging this new dataset, we quantitatively highlight key disparities between AI methods and scientific problems, revealing substantial opportunities for deeper AI integration across scientific disciplines. Furthermore, we explore the potential and challenges of facilitating collaboration between AI and scientific communities through the lens of link prediction. Our findings and tools aim to promote more impactful interdisciplinary collaborations and accelerate scientific discovery through deeper and broader AI integration. Our code and dataset are available at: https://github.com/charles-pyj/Bridging-AI-and-Science.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Applied computingccs: Information systems Web miningccs: Information systems Collaborative and social computing systems and tools
@@ 1. Introduction
The 2024 Nobel Prizes in Physics and Chemistry were both awarded to Artificial Intelligence (AI) researchers. In particular, the developers of AlphaFold (Jumper et al., 2021) were recognized for their groundbreaking work, which is based on the Transformer network and has revolutionized the prediction of protein structures, providing a transformative approach to solving complex biological problems. This compellingly exemplifies how AI has emerged as a powerful tool to facilitate scientific research across various disciplines (Wang et al., 2023; Leontidis, 2024; Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
@@
However, despite the transformative potential of using AI methods for solving scientific problems (AI4Science), a considerable gap persists between AI and scientific communities, hindering the full exploitation of AI for scientific discovery: On the one side, advanced AI methodologies might remain mysterious or underutilized by scientists; On the other side, AI researchers may lack awareness of the specific challenges and potential applications in scientific domains, missing opportunities for interdisciplinary collaborations.
These challenges invoke a few key research questions: What does the AI4Science landscape look like? Are AI methods and scientific problems evenly interconnected, or are some of them underexplored? And how could we expand the connections?
@@
Existing efforts to approach these questions have predominantly involved small-scale, qualitative reviews of AI’s application in science (Wang et al., 2023; Leontidis, 2024), particularly in specialized areas such as drug discovery and material science (Mak et al., 2023; Guo et al., 2021; Xu et al., 2024). These qualitative analyses of the literature often rely on heuristic insights from domain experts to suggest potential uses of AI in solving scientific problems. While valuable, such focused reviews are limited in providing a comprehensive and diverse perspective on the AI4Science landscape.
@@
In the mean time, recent research from the Science of Science community has initialized larger-scale studies that quantify AI’s impact on publications in scientific fields (Gao and Wang, 2023; Duede et al., 2024), although they often heavily rely on established scientific taxonomies in “AI-heavy” fields. There remains a lack of delivering a holistic, dynamic, and data-driven overview of AI4Science. Such analysis is crucial for understanding the barriers and identifying new opportunities for deep engagement of fast-evolving AI methodology in scientific research.
@@
Aiming at addressing these key questions, we conduct a large-scale and comprehensive analysis of relevant literature over the past decade. We start by using large language models (LLMs) to identify the scientific problems and AI methods addressed in publications from top science and AI venues, through which we assemble a novel and balanced AI4Science dataset to analyze the role of AI in scientific research (Sec. 2). Leveraging this dataset, we illustrate the AI4Science landscape through two projection maps (Fig. 2) and a bipartite graph (Fig. 4a). We further quantitatively reveal the current discrepancies between AI and science, leading to several key findings and implications (Sec. 3). Finally, through the lens of link prediction, we demonstrate the potential and challenges in advancing the bridge between AI and science (Sec. 4). Our findings and tools may foster more effective interdisciplinary collaborations and accelerate scientific discovery through a more diverse and deeper integration of AI methods.
Figure 1. Illustration of LLM-based extraction of scientific problem $p_{i}$, AI method $m_{i}$, and AI usage $u_{i}$ from an example publication. The extractions are then semantically projected, clustered, and labeled to form the AI4Science landscape (Fig. 2), as well as to construct the bipartite graph (Fig. 4a).
2. A New Dataset and the AI4Science Landscape
@@
A major challenge in understanding the disparities and potential of AI4Science is the lack of a comprehensive and balanced dataset. Existing literature data are often biased toward successful stories, such as areas where AI and science are already deeply intertwined. Or they reflect the perspective of either the scientific or the AI communities, but not both. We take the initiative to create a broad and balanced large-scale literature dataset that provides a wide-angled lens for understanding AI4Science research.
@@ 2.1. Curation of AI4Science Literature
  • For science domains, we include three top multidisciplinary journals that represent cutting-edge and high-quality research (Nature, Science, and PNAS) and two of their subjournals (Nature Communications and Science Advances).
  • For AI communities, we include seven top conferences from the list of CSRankings.org, comprising two AI-focused venues (AAAI, IJCAI), three machine learning venues (ICLR, ICML, NeurIPS), and two venues of applied AI research (SIGKDD, WWW).
Our data includes a total of 162,656 publications from 2014 to 2024. Appendix A.1 provides details of the data curation process.
Figure 2. 2D projection maps of the AI4Science landscape: (a) the extracted scientific problems $\{p_{i}\}$ and (b) AI methods $\{m_{i}\}$. Each dot represents a publication, with coordinates obtained by reducing the dimensionality of the problem/method embeddings. Green dots correspond to AI4Science work, using AI methods to address scientific problems; Orange dots show papers addressing scientific problems without using AI methods; Purple dots are papers using AI to address non-scientific problems. The visualizations reveal a noticeable discrepancy in the distribution of AI4Science work (green) versus non-AI4Science work (orange and purple) in both the problem and method spaces.
@@ 2.2. Extracting Scientific Problems and AI Methods with LLMs
To understand how AI is applied to science, we need to extract paired entries of scientific problems and AI solutions from relevant publications. Traditionally, this is accomplished through literature reviews and surveys which involve manual examinations of small samples of papers (Wang et al., 2023; Leontidis, 2024), a process that is labor-intensive and time-consuming. Recent quantitative studies in the Science of Science field have started using metadata entries, such as keywords, to identify AI usage and categorize scientific topics (Gao and Wang, 2023; Duede et al., 2024). While this method enables larger-scale analysis, it offers only a coarse-grained view, potentially missing the nuanced details of scientific problems and AI methods. Additionally, it can suffer from issues such as incomplete data entries, limiting the effectiveness of analysis.
@@
It is possible that a publication did not address a scientific problem or did not use an AI method, in which case the corresponding field(s) would remain empty. If a paper addresses a scientific problem using an AI method – i.e., $p_{i}$, $m_{i}$, and $u_{i}$ are all non-empty – it is considered as an AI4Science work. Fig. 1 presents an example of the extraction results of an AI4Science publication.
@@
2.3. Semantic Projection and Clustering
One of the key advantages of this new AI4Science dataset compared to the data used in existing studies (Gao and Wang, 2023; Duede et al., 2024) is that our dataset contains detailed textual descriptions of the identified scientific problems and AI methods rather than pre-defined taxonomies, allowing for a deeper semantic analysis of their relationships.
To achieve this, we generate semantic embeddings for the scientific problems $\{p_{i}\}$ and AI methods $\{m_{i}\}$ using the InstructorEmbedding model (Su et al., 2022). We then apply LargeVis (Tang et al., 2016) to project these high-dimensional embeddings into 2D coordinates. Next, a density-based clustering method, HDBSCAN (Campello et al., 2013) is utilized to group similar scientific problems and AI methods in this 2D space. For each cluster, we use GPT-4o to generate an understandable summarization based on frequent keywords and samples of extraction results. This process results in 390 and 355 clusters respectively, with corresponding cluster labels $P_{i}$ and $M_{i}$ to each data point. The choice of the methods and the details of the process are in Appendix A.3.
Fig. 2 shows the 2D maps that illustrate the AI4Science landscape. Fig. 5-6 in Appendix present the maps with the cluster labels.
Unlike using keywords or categorizations from publication metadata (Gao and Wang, 2023; Duede et al., 2024), our approach employs a data-driven methodology. By using textual embedding, clustering, and topic labeling, relevant AI methods discussed in both scientific and AI literature are grouped together, as are the scientific problems mentioned in both fields. This enables a semantic understanding of the AI4Science space and ensures greater consistency and generalizability of the data.
@@ 2.4. Dataset Overview and the Bipartite Graph
To better illustrate the connections between AI and science, we construct a bipartite graph based on the clusters. In this bipartite graph, scientific problem clusters and AI method clusters are represented as two types of nodes, and the publications act as edges connecting them. Formally, the graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{P_{i}\}\cup\{M_{i}\}$ and $\mathcal{E}=\mathcal{D}$. As shown in Fig. 4a, the graph provides an intuitive visual representation of how AI is linked to scientific challenges and vice versa.
Notation Value
Publication year range 2014–2024
# of venues 12
# of publications $N$ 162,656
# of scientific problem extractions |$|\{p_{i}\}|$ 129,038
# of scientific problem clusters |$|\{P_{i}\}|$ 390
# of AI method extractions |$|\{m_{i}\}|$ 42,455
# of AI method clusters |$|\{M_{i}\}|$ 355
# of AI usage extractions |$|\{u_{i}\}|$ 42,455
# of AI4Science publications 7,542
@@
Figure 3. Scientific problem clusters (a) and AI method clusters (b) are visualized as scatters. In both plots, the $x$-axis represents the total number of publications in each cluster, while the $y$-axis reflects the number of interdisciplinary AI4Science publications in each cluster. The black lines show the regression results on the clusters. Clusters above the line indicate regions well-explored for AI4Science. Clusters falling below the line highlight the under-explored regions, where the integration of AI and science remains limited. These areas represent potential opportunities for further interdisciplinary collaboration.
@@ 3. Discrepancy Between AI and Science
Both Fig. 2 and 4a show a noticeable pattern of discrepancy between AI and science. This section quantitatively assesses these disparities, providing a deeper understanding of the gaps.
@@ 3.1. Uneven Distribution of AI4Science Research
The projection maps in Fig. 2 illustrate the distribution of AI4Science work within the semantic spaces of scientific problems and AI methods (the green dots among orange and purple dots). The uneven distributions suggest that AI4Science work is heavily clustered in certain subareas, while it remains underrepresented in wide ranges of both the problem and the method spaces.
In Fig. 3, we plot the clusters of scientific problems and AI methods, showing the relation between cluster sizes and the presence of AI4Science publications within each cluster. The slopes of the regression lines represent the average proportions of AI4Science work in the literature of each scientific problem cluster or AI method cluster. Clusters below the regression line indicate problems or AI methods in which the integration of AI4Science is underrepresented (noted the under-explored regions), suggesting potential opportunities for further interdisciplinary collaboration.
11footnotetext: Six outstanding data points are skipped from the scatter plots for visualization clarity: ‘Neural Network Challenges’ (820,375) and ‘Urban Traffic Management’ (551,298) in (a), ‘Machine Learning’ (770,539), ‘Deep Learning Models’ (605,344), ‘Genomic Analysis Methods’ (292,287), and ‘Molecular Dynamics Simulations’ (214,210) in (b).
Underexplored scientific problems.
In the bottom-right section of Fig. 3a, below the regression line, several scientific problem clusters are highlighted as being rarely approached through AI techniques. Notable examples include ‘Asymmetric Synthesis’, ‘Magnetic Materials and Spintronics’, ‘Forests and Climate Change’, ‘Viral Mechanisms’, ‘Evolutionary Paleontology’, ‘Nanostructure Synthesis’, ‘Diabetes and Insulin Regulation’, ‘Bacterial Infections and Immune Response’, ‘Cancer Signaling Pathways’, and ‘Cancer Drug Resistance’. Scientific problems in this region could benefit from more engagement of AI methods.
Underutilized AI methods.
Similarly, Fig. 3b reveals AI techniques that have not yet been widely applied to sciences, including: ‘Attention Mechanisms’, ‘Gradient-Based Methods’, ‘Graph Embedding’, ‘Regularization Methods’, ‘Transformer Architectures’, and ‘Contrastive Learning’. These techniques could have been utilized in a wider range of scientific applications.
Figure 4. The AI-Science bipartite graph and node distributions. (a) The bipartite graph with scientific problem clusters and AI method clusters as nodes. The size of each node corresponds to its unweighted degree, representing the number of AI methods applied to a scientific problem, or the number of scientific problems utilizing an AI method. For visualization clarity, edges representing fewer than four publications and the resulting isolated noes are hidden. The distribution of links indicates imbalanced connectivity. (b-c) The degree of AI method nodes follows a log-normal distribution, and the degree distribution of scientific problem nodes is even more heavily tailed, suggesting the existence of “hubs” in linking AI and science.